for a group of men of similar ages in all occu- pations Questions 11, 12, and 13 are based on the infor mation given below Numbers of People and Deaths from Disease Z by Age Group in Communities X and Y COMMUNITY X COMMUNITYY No. of Deaths No. of Deaths Age Group No. of No. of From Dis From Dis People People ease Z ease Z 69 48 Young 8.000 5.000 Old 11,000 115 3.000 Calculate the age-adjusted death rate for dis- ease Z in communities X and Y by the direct method, using the total of both communities as the standard population 11. The age-adjusted death rate from disease Z for community X is: 12. The proportionate mortality from disease Z for community Y is a. 9.6/1,000 b. 13.5/1,000 .20.0/1,000 C
6. For a disease such as pancreatic cancer which is highly fatal and of short duration: a. Incidence rates and mortality rates will be similar b. Mortality rates will be much higher than in- cidence rates c. Incidence rates will be much higher than mortality rates d. Incidence rates will be unrelated to mortal ity rates None of the above
Answer #1

The total number of population is 27000 for the bothe community X and Y .

11. Age adjusted death rate from disease Z for community X is :

Number of young people = 69

Number of old people = 115

Total population = 19000

Death rate = 69+ 115 /× 100

= 184 / 19000×100

= 0.96%

12. Proportionate mortality from disease Z :-

Total number of population in community Y = 8000

Number of young people = 48

Number of old people = 60

= 48+60 /8000 × 100

= 108/8000 ×100

= 13.5 %

6) The correct option is b) Mortality rate will be much higher than incidence rates .